How do you simplify (3-2radical 12)squared?
Answer by RAY100(1637)

(3-2sqrt12)^2
(3-2sqrt12) (3-2sqrt12)
9-12sqrt12+4*12
9+48-12sqrt(4*3)
57-12*2sqrt3
57-24sqrt3
check
57-24sqrt3=15.43
(3-2sqrt12)^2=15.43 ok
things to know (x+1)^2=(x+1)(x+1)
(sqrt 3)^2=3
or (sqrt9)^2= 3^2=9
sqrt(300)=sqrt(3*100)=sqrt3*sqrt100=10sqrt3